OpenStudy (anonymous):

Vectors Just to double-check whether my working is correct. @ganeshie8

OpenStudy (anonymous):

\[Given~u=\left(\begin{matrix}6 \\ 2\end{matrix}\right),v=\left(\begin{matrix}h \\ -1\end{matrix}\right)\]Find the possible values of h for each of the following cases: a)u and v are parallel

ganeshie8 (ganeshie8):

hold up

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

do you know how to draw the vector \(u=\left(\begin{matrix}6 \\ 2\end{matrix}\right)\) on xy plane ?

ganeshie8 (ganeshie8):